# Superdiffusion in random two dimensional system with time-reversal symmetry and long-range hopping

### v3

@@ V.3 Transport
Finally, we consider the particle transport that is the main distinction of two phases. As we explained earlier it is expected to be superdiffusive in the fast phase and diffusive in the slow phase. For the initially localized particle its typical displacement $R$ increases with the time following a sort of diffusion law $R^{2}\propto C(R)t$ since conductance and diffusion coefficient are synonyms in two dimensions. In the fast phase a conductance $C(R)$ increases logarithmically with the size $R$ leading to a superdiffusive behavior in contrast to the linear dependence expected in the slow phase where conductance remains finite. This expectation is consistent with Ref. [11], where the behavior $R\propto t^{1/d}$ was predicted for the slow phase in a $d$- dimensional system. For $d=2$ this is equivalent to the diffusive behavior in contrast to the subdiffusive behavior for $d=3$ [11] or the superdiffusive behavior for $d=1$ [12].
@@
We do not consider the most often used root mean square displacement because the power-law tails of the wavefunction can lead to the overestimate of the actual move. Indeed, the wavefunction asymptotic behavior $\psi(R)\propto R^{-2}$ emerges in the first order perturbation theory approach to the present problem if hopping is treated as a perturbation. Although such behavior suggests the particle localization nearby the origin $R=0$, the average squared displacement for that asymptotic behavior diverges logarithmically with the system size, while logarithmic averaging does not lead to any divergence.
The results of the calculations for $R_{\rm log}(t)^{2}$ vs. $t$ are shown in Fig. 4. Based on these results it is still difficult to characterize the transport because at very short times $t<1$, $1/W$ one has $|c_{i}(t)|\approx V_{i}t$ ($V_{i}$ is the coupling strength of the initial site and the site $i$) leading to $R_{\rm log}(t)\propto t$ , while the saturation at around the system size takes place at long times $t>100$. To focus on the relevant time domain for the superdiffusive transport we restrict our consideration to times $t_{min}<t<t_{max}$ with $t_{min}=1$ and $t_{max}=100$. The time dependence of the relative conductance $C_{*}(t)/C_{*}(1)$ ($C_{*}(t)=R_{\rm log}(t)^{2}/t$ in that time domain in shown in Fig. 5 for various disorder strengths. Relative conductance is used since we are interested in the time dependence of the diffusion coefficient (conductance) rather than its absolute value. If the transport is superdiffusive this effective conductance should increase with the time. logarithmically, saturating at long times due to a finite size effects.
According to Fig. 5 the superdiffusive transport, indeed, emerges for the weakest disorder at $W\leq 1.5$ in accord with the theoretical expectations. At short times the conductance time dependence is consistent with the expected logarithmic growing for $W=0.1$ and $W=0.5$, while for $W=1.5$ the growing domain is too narrow to make any conclusions possibly due to finite size effects. For stronger disorder the conductance slowly decreases with the time. At the transition point $W\approx 2.5$ corresponding to the zero energy, Eq. (17) predicts the superdiffusive behavior $C\propto\sqrt{\ln(L)}$, that we do not see, possibly because of dominating contributions of slow states with energies different from 00. The reduction of the diffusion coefficient (conductance) with the time (size) for $W>1.5$ can be due to the renormalization of coupling constant occurring for the dipolar hopping and dipoles oriented within the same direction for strong disorder [71]. Although disorder is not strong, some reduction of the effective coupling constant can still take place.
@@ Appendix A Fourier transforms of hopping amplitudes
The anisotropic dipole-dipole interaction, responsible for the hopping in the model under consideration, is defined by Eq. (4). Here we calculate numerically its Fourier transform $V(\mathbf{q})=\sum_{j}V_{ij}e^{i\mathbf{qr}_{ij}}$ needed to evaluate the zeroth order conductance, Eq. (8).
@@
$V(\mathbf{q})=V_{0}\pi\frac{-q_{x}^{2}+q_{y}^{2}}{q^{2}},$ (29)
This approximation works reasonably well for the periodic square lattice of the size $L=100$ as illustrated in Fig. 6. It becomes exact for the system size approaching infinity since the logarithmic divergence of the zeroth order conductance emerges at $q\rightarrow 0$ corresponding to long distances. In this limit the regular correction to Eq. (29) disappears because the sum of all dipole-dipole interactions is zero.
@@
Figure 7: Comparison of the numerical estimate of eigenstate fractal dimensions with the analytical theory of Ref. [28]. The only exception is the case of $W=3$ for the present graph, where we cannot fit those data even choosing the minimum possible value of $L_{0}\rightarrow 0$, corresponding to the infinite length limit of the fractal dimension., as shown by the dashed line..
@@
In Figs. 7, 8 we show the comparison of the numerical results for the dimension $D_{1}$ (the numbers of realizations are as in Fig. 3) and the analytical theory $2-c_{loc}/C$ for two choices of the parameter $\zeta=1/2$ and $1$, respectively. It is clear from Figs. 7, 8 that in both cases the analytical theory does not provide an acceptable fit of the data. However, if we set $D_{1}=2-1.3c_{loc}/C$, and employ Eq. (30) with the parameter $\zeta=1$, then we get an almost perfect agreement of numerical and analytical results as reported in the main text, Fig. 3.

### v4

@@ V.3 Transport
Finally, we consider the particle transport that is the main distinction of two phases. Our goal here is just to demonstrate that the superdiffusive transport, indeed, exists at small disorder strength in a reasonable agreement with the predictions of the analytical theory. We do not attempt to compare the predictions of analytical theory with numerical results in detail because the transport includes many eigenstates with different energies and different transport rates that makes an accurate analytical consideration overcomplicated.
As we explained earlier it is expected to be superdiffusive in the fast phase and diffusive in the slow phase. For the initially localized particle its typical displacement $R$ increases with the time following a sort of diffusion law $R^{2}\propto C(R)t$ since conductance and diffusion coefficient are synonyms in two dimensions. In the fast phase a conductance $C(R)$ increases logarithmically with the size $R$ leading to a superdiffusive behavior in contrast to the linear dependence expected in the slow phase where conductance remains finite. This expectation is consistent with Ref. [11], where the behavior $R\propto t^{1/d}$ was predicted for the slow phase in a $d$- dimensional system. For $d=2$ this is equivalent to the diffusive behavior in contrast to the subdiffusive behavior for $d=3$ [11] or the superdiffusive behavior for $d=1$ [12].
@@
We do not consider the most often used mean squared displacement because the power-law tails of the wavefunction can lead to the overestimate of the actual move. Indeed, even at short times a mean squared displacement diverges with the system size because of the wavefunction asymptotic behavior $\psi(R)\propto R^{-2}$ emerging in the first order perturbation theory in a hopping. Although at short times the particle is localized nearby the origin $R=0$, the average squared displacement $\int d^{2}\mathbf{R}|\psi(R)|^{2}R^{2}$ diverges logarithmically for $\psi(R)\propto R^{-2}$ for the system of an infinite size, while logarithmic averaging does not lead to any divergence. at a finite time.
The results of the calculations for $R_{\rm log}(t)^{2}$ vs. $t$ are shown in Fig. 5. Based on these results it is still difficult to characterize the transport because at very short times $t<1$, $1/W$ one has $|c_{i}(t)|\approx V_{i}t$ ($V_{i}$ is the coupling strength of the initial site and the site $i$) leading to $R_{\rm log}(t)\propto t$ , while the saturation at around the system size takes place at long times $t>100$. To focus on the relevant time domain for the superdiffusive transport we restrict our consideration to times $t_{min}<t<t_{max}$ with $t_{min}=1$ and $t_{max}=100$. The time dependence of the relative conductance $C_{*}(t)/C_{*}(1)$ ($C_{*}(t)=R_{\rm log}(t)^{2}/t$ in that time domain in shown in Fig. 6 for various disorder strengths. Relative conductance is used since we are interested in the time dependence of the diffusion coefficient (conductance) rather than its absolute value. If the transport is superdiffusive this effective conductance should increase with the time. logarithmically, saturating at long times due to a finite-size effects.
According to Fig. 6 the superdiffusive transport, indeed, emerges for the weakest disorder at $W\leq 1.5$ in accord with the theoretical expectations. At short times the conductance time dependence is consistent with the expected logarithmic growing for $W=0.1$ and $W=0.5$, while for $W=1.5$ the growing domain is too narrow to make any conclusions possibly due to finite-size effects. For stronger disorder the conductance slowly decreases with the time. At the transition point $W\approx 2.5$ corresponding to the zero energy, Eq. (18) predicts the superdiffusive behavior $C\propto\sqrt{\ln(L)}$, that we do not see, possibly because of dominating contributions of slow states with energies different from 00. The reduction of the diffusion coefficient (conductance) with the time (size) for $W>1.5$ can be due to the renormalization of coupling constant occurring for the dipolar hopping and dipoles oriented within the same direction for strong disorder [74]. Although disorder is not strong, some reduction of the effective coupling constant can still take place.
@@ Appendix A Fourier transforms of hopping amplitudes
The anisotropic dipole-dipole interaction, responsible for the hopping in the model under consideration, is defined by Eq. (4). Here we calculate numerically its Fourier transform $V(\mathbf{q})=\sum_{j}V_{ij}e^{i\mathbf{qr}_{ij}}$ needed to evaluate the classical conductance, Eq. (9).
@@
$V(\mathbf{q})=2V_{0}\pi\frac{q_{y}^{2}-q_{x}^{2}}{q^{2}},$ (27)
This approximation works reasonably well for the periodic square lattice of the size $L=100$ as illustrated in Fig. 7. It becomes exact for the system size approaching infinity since the logarithmic divergence of the classical conductance emerges at $q\rightarrow 0$ corresponding to long distances. In this limit the regular correction to Eq. (27) disappears because the sum of all dipole-dipole interactions is zero.
@@
Figure 8: Comparison of the numerical estimate of eigenstate fractal dimensions with the analytical theory of Ref. [28]. The only exception is the case of $W=3$ for the present graph, where we cannot fit those data even choosing the minimum possible value of $L_{0}\rightarrow 0$, corresponding to the infinite-length limit of the fractal dimension., as shown by the dashed line..
@@
In Figs. 8, 9 we show the comparison of the numerical results for the dimension $D_{1}$ (the numbers of realizations are as in Fig. 4) and the analytical theory $2-c_{loc}/C$ for two choices of the parameter $\zeta=1/2$ and $1$, respectively. It is clear from Figs. 8, 9 that in both cases the analytical theory does not provide an acceptable fit of the data. However, if we set $D_{1}=2-1.3c_{loc}/C$, and employ Eq. (28) with the parameter $\zeta=1$, then we get an almost perfect agreement of numerical and analytical results as reported in the main text, Fig. 4.